OpenStudy (anonymous):

determine the solution of the system. y=x+6 and y=2x

OpenStudy (anonymous):

y=x+6....1 y=2x.......2 put eqn.1 in eqn.2 x+6=2x 2x-x=6 x=6 put the value of x in any of the two equations. y=2(6) y=12

OpenStudy (anonymous):

I have another one for u can u help I have a couple that I am stuck on

OpenStudy (anonymous):

yh